if I have a seven sided i guess you would call it irregular and it wants you to find X and the interior angles are marked as the following 1=125deg ,2=135deg,3=130deg,4=129deg,5=116deg,6=120deg 7=x___deg
i used the numbers 1-7 to just number the interior angles

Answer by ikleyn(52863)

.
Use the fact that the sum of interior angles of the 7-gon is equal to (7-2)*180 = 5*180 = 900 degrees.


Therefore, to find the unknown 7-th angle, distract the sum of the given six angles from 900 degrees.